OpenStudy (anonymous):

Ehat is the volume of a sphere with radius 9 ft? Leave answers in terms of pi please.

hartnn (hartnn):

volume of sphere with radius r is \(\huge V=(4/3)\pi r^3\) just put r = 9 here.

OpenStudy (anonymous):

Is it 108 pi ft to the third power?

hartnn (hartnn):

nopes, r^3 = 9^3 = 729 4/3 *729 = 972 so answer is 972 pi ft^3

OpenStudy (anonymous):

Do you know what I did wrong?

hartnn (hartnn):

how did u get 108 ?